what is the concentration of potassium dichromate in a solution prepared as 20mg/2ml of 6M sulfuric acid using 0.1015g of potassium dichromate
Answer #1

Molar mass of potassium dichromate = 294 g/mol

moles of potassium dichromate = 0.1015/294 = 3.45*10-4

Volume of the solution = 2 ml = 2/1000 = 0.002 L

Molarity = moles of solute / Volume of solution ( L)

molar concentration of potassium dichromate = 3.45*10-4 / volume (L)

= 3.45*10-4/0.002

= 0.1725 M